OpenStudy (anonymous):

\[\int\limits_{-6}^{-3}{1 \over x+2}dx\] Please help with the steps of integration.. I think I can put the -3 and -6 in the equation after you help me integrating it.

OpenStudy (campbell_st):

this is a ln function since the numerator is the derivative of the denominator f(x) = x + 2 f'(x) = 1 so \[\int\limits_{-6}^{-1} 1/(x+2) dx = \ln(x + 2)\] so F(x) = ln(x + 2) Now evaluate by F(-3) - F(-6) so ln(-1) - ln(-4)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Denyekwe, if you're still there, can you check my new post?

OpenStudy (anonymous):

Look at this image and let me know what you think

OpenStudy (anonymous):

It makes quite some sense :)

OpenStudy (anonymous):

What is confusing about it

OpenStudy (anonymous):

It's not confusing, but does it work for every thing like that?

OpenStudy (anonymous):

Yes as long as it is in that form. Notice that it is just X not (x^something)

OpenStudy (anonymous):

Yes. Thank you!